# Kudla Millson lift of toric cycles and restriction of Hilbert modular forms

### v3

@@ 3.4.  Restriction of scalars and seesaw
In section 3.1 we introduced the notations $\mathcal{D}$, $Y$ and $\mathcal{C}$ attached to the quadratic space $\mathcal{V}$ where $\mathcal{V}$ is either the quadratic space $V$ over $\mathbb{Q}$ or the quadratic space $E$ over $F$. In the later case, we will replace the notation by $\mathcal{D}_{T}$, $Y_{T}$ , $\mathcal{C}_{T}$ and $K_{T}$, as used in the introduction.
@@
$\omega_{\varphi}\colon\operatorname{SL}_{2}(\mathbb{A})\times\operatorname{SO}_{V}({\mathbb{A}})\longrightarrow\mathcal{S}(V_{\mathbb{A}}).$ (3.75)
Let $\varphi_{f}=\mathbf{1}_{\widehat{L}}\in\mathcal{S}(V_{\widehat{\mathbb{Q}}})$ be as previously, the characteristic function of an even integral lattice of level $N$. Let $\varphi=\varphi_{\operatorname{KM}}^{V}\otimes\varphi_{f}$, where $\varphi_{\operatorname{KM}}^{V}\in\Omega^{q}(\mathcal{D}^{+})$ is the Kudla-Millson form on $\mathcal{D}^{+}$. In this setting, the Kudla-Millson lift is
$\Theta_{\varphi}\colon H_{q}(Y)\longrightarrow M_{d}(\Gamma_{0}(N)).$ (3.76)
@@
We can view the Schwartz function $\varphi_{f}\in\mathcal{S}(V_{\widehat{\mathbb{Q}}})$ used for the lift (3.76) as a finite Schwartz function in $\varphi_{f}\in\mathcal{S}(\widehat{E})$ and set $\varphi=\varphi_{\operatorname{KM}}^{E}\otimes\varphi_{f}$, where $\varphi_{\operatorname{KM}}^{E}\in\Omega^{q}(\mathcal{D}_{T}^{+})\simeq\Omega^{q}(\mathbb{R}_{>0}^{q})$ is the Kudla-Millson form on $\mathcal{D}_{T}^{+}$.
Let $K_{T,f}\times K_{T,f}^{\prime}\subset\operatorname{SO}_{E}(\widehat{F})\times\operatorname{SL}_{2}(\widehat{F})$ be the open compacts stabilizing $\varphi$ under the Weil representation, and let $Y_{T}$ the locally symmetric space. Let $\mathcal{C}_{T}$ be the finite group of double cosets
$\mathcal{C}_{T}\coloneqq\operatorname{T}(\mathbb{Q})^{+}\backslash\operatorname{T}(\widehat{\mathbb{Q}})/K_{T,f}=E^{1,+}\backslash\widehat{E}^{1}/K_{T,f}$ (3.79)
@@
$Y_{T}=\operatorname{T}(\mathbb{Q})\backslash\operatorname{T}(\mathbb{A})/K_{T}\simeq E^{1}\backslash E^{1}_{\mathbb{A}}/K_{T}\simeq E^{1}\backslash\left((\mathbb{R}_{>0})^{q}\times\widehat{E}^{1}\right)/K_{T,f}$ (3.80)
@@
where $\Lambda\coloneqq\operatorname{T}(\mathbb{Q})^{+}\cap K_{T,f}$. By [PR94, Theorem. 4.11, p. 208], the quotient $Y_{T}$ is compact since $\operatorname{T}(\mathbb{Q})\simeq E^{1}$ is $\mathbb{Q}$-anisotropic. Then $\dim(Y_{T})={q}$ and the Kudla-Millson theta lift is
$\Theta_{\varphi}\colon H_{q}(Y_{T})\longrightarrow M_{(1,\dots,1)}(\Gamma^{\prime}),$ (3.82)
where $\Gamma^{\prime}$ is a congruence subgroup of $\operatorname{SL}_{2}(F)$.
We have $H_{q}(Y_{T},\mathbb{C})\simeq\mathbb{C}[\mathcal{C}_{T}]$. Let $Y_{T,I}$ be the connected component corresponding to $I$ in (3.81) and $[Y_{T,I}]\in H_{q}(Y_{T,I})$ a fundamental class, i.e. a generator of $H_{q}(Y_{T,I},\mathbb{Z})\simeq\mathbb{Z}$. Let $\chi_{f}$ be the finite part of the Hecke character and suppose that $\chi_{f}$ is trivial on $K_{T,f}$. Hence, we can view $\chi_{f}$ as a function on $\mathcal{C}_{T}$ and define
@@
$\Theta_{\varphi}(\tau_{1},\cdots,\tau_{d},\chi)\coloneqq\int_{Y_{\chi}}\Theta_{\varphi}(\tau_{1},\cdots,\tau_{d})\in M_{(1,\dots,1)}(\Gamma^{\prime}).$ (3.84)
@@
$\Theta_{\varphi}(\tau_{1},\cdots,\tau_{d})=\widetilde{\Theta}_{\varphi}(\tau_{1},\cdots,\tau_{d},t)dt^{\times}\in\Omega^{q}(Y_{T})$ (3.86)
@@
$\operatorname{SL}_{2}(F)$⟶$\lhook\joinrel\longrightarrow\operatorname{Sp}_{4d}(\mathbb{Q})⸦$
@@
[…]
We also had the pairs of open compact subgroups $K_{f}\times K_{f}^{\prime}\subset\operatorname{SO}_{V}(\widehat{\mathbb{Q}})\times\operatorname{SL}_{2}(\widehat{\mathbb{Q}})$ and $K_{T,f}\times K_{T,f}^{\prime}\subset\operatorname{SO}_{E}(\widehat{F})\times\operatorname{SL}_{2}(\widehat{F})$ stabilizing $\varphi_{f}$. Suppose they are compatible with the embeddings in the sense that $K_{T,f}\subset K_{f}\cap\operatorname{SO}_{E}(\widehat{F})$ and $K_{f}^{\prime}\subset K_{T,f}^{\prime}\cap\operatorname{SL}_{2}(\widehat{\mathbb{Q}})$. After conjugating if necessary, we can assume that the maximal compacts $K_{\infty}$ and $K_{T,\infty}$ are chosen such that $K_{\infty}\cap\operatorname{SO}_{V}(\mathbb{R})^{+}=K_{T,\infty}$. Then the embedding of $\operatorname{SO}_{E}$ in $\operatorname{SO}_{V}$ induces an immersion
@@
On the other hand, the pullback by the diagonal inclusion of $\mathscr{H}$ in $\mathscr{H}\times\cdots\times\mathscr{H}$ induces the diagonal restriction
$\iota^{\ast}\colon M_{(1,\dots,1)}(\Gamma^{\prime})\longrightarrow M_{d}(\Gamma_{0}(N)).$ (3.96)
The seesaw identity relies on two observations. Firstly, the two Weil representations $\omega_{V}$ and $\omega_{E}$ agree on their smallest common subgroups of both dual pairs, namely $(\operatorname{SO}_{E}(F),\operatorname{SL}_{2}(\mathbb{Q}))$. Secondly, the Kudla-Millson form has the functorial property that ${\left.\kern-1.2pt\varphi_{\operatorname{KM}}^{V}\vphantom{\big{|}}\right|_{\mathcal{D}_{T}^{+}}}=\varphi_{\operatorname{KM}}^{E}$, see [Bra23a, Equation. (5.10)]. Hence, the two kernels
$\Theta_{\varphi}\in C^{\infty}\left(\operatorname{SL}_{2}(\mathbb{Q})\backslash\operatorname{SL}_{2}(\mathbb{A})\right)\otimes\Omega^{q}(Y)$ (3.97)
@@
$\Theta_{\varphi}\in C^{\infty}\left(\operatorname{SL}_{2}(F)\backslash\operatorname{SL}_{2}(\mathbb{A}_{F})\right)\otimes\Omega^{q}(Y_{T})$ (3.98)
@@
$C^{\infty}\left(\operatorname{SL}_{2}(\mathbb{Q})\backslash\operatorname{SL}_{2}(\mathbb{A})\right)\otimes\Omega^{q}(Y_{T}).$ (3.99)
@@
$\iota^{\ast}\Theta_{\varphi}(\tau_{1},\dots,\tau_{d})=\Theta_{\varphi}(\tau,\dots,\tau)=\phi^{\ast}\Theta_{\varphi}(\tau).$ (3.100)
@@
$\Theta_{\varphi}(\tau,\cdots,\tau,\chi)=\int_{Y_{\chi}}\Theta_{\varphi}(\tau,\cdots,\tau)=\int_{Y_{\chi}}\phi^{\ast}\Theta_{\varphi}(\tau)=\int_{\phi_{\ast}Y_{\chi}}\Theta_{\varphi}(\tau),$ (3.103)
@@
$\Theta_{\varphi}(\tau,\cdots,\tau,\chi)=\int_{C_{\chi}}\Theta_{\varphi}(\tau).$ (3.104)
@@
For now we have only considered the case where $E$ is an extension of a field $F$, but we want to consider more general étale algebras of the form $E=E_{1}\times\cdots\times E_{r}$, together with an involution fixing $F=F_{1}\times\cdots\times F_{r}$. Here each of the $F_{i}$ is a totally real field of degree $d_{i}$, that is the fixed field of $E_{i}$. Suppose that $(V,Q)$ is the restriction of scalars of $(E,Q_{\alpha})$ from $F$ to $\mathbb{Q}$, as in 2.1. Then $V$ is a vector space of dimension $2d$ where $d=d_{1}+\cdots+d_{r}$. Furthermore, we can split the quadratic space as $V=V_{1}\oplus\cdots\oplus V_{r}$, where $V_{i}$ is the restriction of scalars of $E_{i}$ from the field $F_{i}$ to $\mathbb{Q}$. We can restrict the embedding $\operatorname{SO}_{V}(\mathbb{Q})\hookrightarrow\operatorname{Sp}_{4d}(\mathbb{Q})$ given in 3.18 to an embedding $\operatorname{SO}_{V_{1}}(\mathbb{Q})\times\cdots\times\operatorname{SO}_{V_{r}}(\mathbb{Q})\hookrightarrow\operatorname{Sp}_{4d}(\mathbb{Q})$. Its commutator is $\operatorname{SL}_{2}(\mathbb{Q})^{r}$ and we get a seesaw
@@
Let $\varphi_{f}\in\mathcal{S}(V_{\widehat{\mathbb{Q}}})$ be a finite Schwartz function. Suppose that $\varphi_{f}$ is the characteristic function of an even integral lattice of level $N$. Let $\varphi_{V}=\varphi^{V}_{\operatorname{KM}}\otimes\varphi_{f}$ and the associated theta lift
$\Theta_{\varphi}\colon H^{q}(Y,\mathbb{C})\longrightarrow M_{d}(\Gamma_{0}(N)).$ (3.105)
We have $\mathcal{S}(V_{\widehat{\mathbb{Q}}})\simeq\mathcal{S}(\widehat{E})\simeq\mathcal{S}(\widehat{E}_{1})\otimes\cdots\otimes\mathcal{S}(\widehat{E}_{r})$ and so we can write $\varphi_{f}$ as a finite sum
with $\varphi_{f}^{\beta,i}\in\mathcal{S}(\widehat{E}_{i})$. Similarly, the character $\chi$ on $\operatorname{T}(\mathbb{Q})\backslash\operatorname{T}(\mathbb{A})$ can be written as $\chi=\chi_{1}\times\cdots\times\chi_{r}$ where
$\chi_{i}\colon\operatorname{T}_{i}(\mathbb{Q})\backslash\operatorname{T}_{i}(\mathbb{A})\longrightarrow\mathbb{C}^{\times}$ (3.107)
@@
of parallel weight one for a congruence subgroup $\Gamma_{i}^{\prime}$ of $\operatorname{SL}_{2}(\mathcal{O}_{F_{i}})$. We denote by
@@
$\int_{C_{\chi}}\Theta_{\varphi}(\tau)=\sum_{\beta}\Theta_{\varphi^{\beta,1}}(\tau^{\Delta},\chi_{1})\cdots\Theta_{\varphi^{\beta,r}}(\tau^{\Delta},\chi_{r}).$ (3.111)

### v4

@@ 3.4. Restriction of scalars and seesaw
In section 3.1 we introduced the notations $\mathcal{D}$, $Y$ and $\mathcal{C}$ attached to the quadratic space $\mathcal{V}$ where $\mathcal{V}$ is either the quadratic space $V$ over $\mathbb{Q}$ or the quadratic space $E$ over $F$. In the later case, we will replace the notation by $\mathcal{D}_{T}$, $Y_{T}$ , $\mathcal{C}_{T}$ and $K_{T}$, as used in the introduction. We will use the notations $\Theta_{\varphi}^{V}$ and $\Theta_{\varphi}^{T}$ if we need to emphasize the underlying dual pair.
@@
$\omega_{V}\colon\operatorname{SL}_{2}(\mathbb{A})\times\operatorname{SO}_{V}({\mathbb{A}})\longrightarrow\mathcal{S}(V_{\mathbb{A}}).$ (3.38)
Let $\varphi_{f}\in\mathcal{S}(V_{\mathbb{A}_{f}})$ a finite Schwartz function as previously. Let $\varphi=\varphi_{\operatorname{KM}}^{V}\otimes\varphi_{f}$, where $\varphi_{\operatorname{KM}}^{V}\in\Omega^{q}(\mathcal{D}^{+})$ is the Kudla-Millson form on $\mathcal{D}^{+}$. In this setting, the Kudla-Millson lift is
$\Theta_{\varphi}\colon H_{q}(Y)\longrightarrow M_{d}(\Gamma^{\prime}_{\mathbb{Q}}),$ (3.39)
for some $\Gamma^{\prime}_{\mathbb{Q}}\subset\operatorname{SL}_{2}(\mathbb{Q})$ depending on $\varphi_{f}$.
@@
We can view the Schwartz function $\varphi_{f}\in\mathcal{S}(V_{\mathbb{A}_{f}})$ used for the lift (3.39) as a finite Schwartz function in $\varphi_{f}\in\mathcal{S}(\mathbb{A}_{E,f})$ and set $\varphi=\varphi_{\operatorname{KM}}^{E}\otimes\varphi_{f}$, where $\varphi_{\operatorname{KM}}^{E}\in\Omega^{q}(\mathcal{D}_{T}^{+})\simeq\Omega^{q}(\mathbb{R}_{>0}^{q})$ is the Kudla-Millson form on $\mathcal{D}_{T}^{+}$.
Let $K_{T,f}\times K_{T,f}^{\prime}\subset\operatorname{SO}_{E}(\mathbb{A}_{F,f})\times\operatorname{SL}_{2}(\mathbb{A}_{F,f})$ be the open compacts stabilizing $\varphi$ under the Weil representation, and let $Y_{T}$ the locally symmetric space. Let $\mathcal{C}_{T}$ be the finite group of double cosets
[…]
@@
$Y_{T}=\operatorname{T}(\mathbb{Q})\backslash\operatorname{T}(\mathbb{A})/K_{T}\simeq E^{1}\backslash E^{1}_{\mathbb{A}}/K_{T}\simeq E^{1}\backslash\left((\mathbb{R}_{>0})^{q}\times\mathbb{A}_{E,f}^{1}\right)/K_{T,f}$ (3.43)
@@
where $\Lambda\coloneqq\operatorname{T}(\mathbb{Q})^{+}\cap K_{T,f}$. By [21, Theorem. 4.11, p. 208], the quotient $Y_{T}$ is compact since $\operatorname{T}(\mathbb{Q})\simeq E^{1}$ is $\mathbb{Q}$-anisotropic. Then $\dim(Y_{T})={q}$ and the Kudla-Millson theta lift is
$\Theta^{T}_{\varphi}\colon H_{q}(Y_{T})\longrightarrow M_{(1,\dots,1)}(\Gamma^{\prime}_{F}),$ (3.45)
where $\Gamma^{\prime}_{F}\subset\operatorname{SL}_{2}(F)$.
We have $H_{q}(Y_{T},\mathbb{C})\simeq\mathbb{C}[\mathcal{C}_{T}]$. Let $Y_{T,I}$ be the connected component corresponding to $I$ in (3.44) and $Y_{T,I}]\in H_{q}(Y_{T,I})$ a fundamental class, i.e. a generator of $H_{q}(Y_{T,I},\mathbb{Z})\simeq\mathbb{Z}$. Let $\chi_{f}$ be the finite part of the Hecke character and suppose that $\chi_{f}$ is trivial on $K_{T,f}$. Hence, we can view $\chi_{f}$ as a function on $\mathcal{C}_{T}$ and define
@@
$\Theta^{T}_{\varphi}(\tau_{1},\cdots,\tau_{d},\chi)\coloneqq\int_{Y_{\chi}}\Theta^{T}_{\varphi}(\tau_{1},\cdots,\tau_{d})\in M_{(1,\dots,1)}(\Gamma^{\prime}_{F}).$ (3.47)
@@
$\Theta^{T}_{\varphi}(\tau_{1},\cdots,\tau_{d})=\widetilde{\Theta}^{T}_{\varphi}(\tau_{1},\cdots,\tau_{d},t)dt^{\times}\in\Omega^{q}(Y_{T})$ (3.49)
@@
$\operatorname{SL}_{2}(F)$ $\lhook\joinrel\longrightarrow\operatorname{Sp}_{4d}(\mathbb{Q})$
@@
$\operatorname{SO}_{V}(\mathbb{Q})$$\operatorname{SL}_{2}(F)$$\operatorname{SO}_{E}(F)$$\operatorname{SL}_{2}(\mathbb{Q}).$}
We also had the pairs of open compact subgroups $K_{f}\times K_{f}^{\prime}\subset\operatorname{SO}_{V}(\mathbb{A}_{f})\times\operatorname{SL}_{2}(\mathbb{A}_{f})$ and $K_{T,f}\times K_{T,f}^{\prime}\subset\operatorname{SO}_{E}(\mathbb{A}_{F,f})\times\operatorname{SL}_{2}(\mathbb{A}_{F,f})$ stabilizing $\varphi_{f}$. Suppose they are compatible with the embeddings in the sense that $K_{T,f}\subset K_{f}\cap\operatorname{SO}_{E}(\mathbb{A}_{F,f})$ and $K_{f}^{\prime}\subset K_{T,f}^{\prime}\cap\operatorname{SL}_{2}(\mathbb{A}_{f})$. After conjugating if necessary, we can assume that the maximal compacts $K_{\infty}$ and $K_{T,\infty}$ are chosen such that $K_{\infty}\cap\operatorname{SO}_{V}(\mathbb{R})^{+}=K_{T,\infty}$. Then the embedding of $\operatorname{SO}_{E}$ in $\operatorname{SO}_{V}$ induces an immersion
@@
On the other hand, the pullback by the diagonal inclusion of $\mathscr{H}$ in $\mathscr{H}^{d}=\mathscr{H}\times\cdots\times\mathscr{H}$ induces the diagonal restriction
$\iota^{\ast}\colon M_{(1,\dots,1)}(\Gamma^{\prime}_{F})\longrightarrow M_{d}(\Gamma^{\prime}_{F}\cap\operatorname{SL}_{2}(\mathbb{Q})).$ (3.55)
Note that $\Gamma^{\prime}_{\mathbb{Q}}\subset\Gamma^{\prime}_{F}\cap\operatorname{SL}_{2}(\mathbb{Q})$ by the assumption on the compact groups $K_{f}^{\prime}$ and $K_{T,f}^{\prime}$. The seesaw identity relies on two observations. Firstly, the two Weil representations $\omega_{V}$ and $\omega_{E}$ agree on their smallest common subgroups of both dual pairs, namely $(\operatorname{SO}_{E}(F),\operatorname{SL}_{2}(\mathbb{Q}))$. Secondly, the Kudla-Millson form has the functorial property that ${\left.\kern-1.2pt\varphi_{\operatorname{KM}}^{V}\vphantom{\big|}\right|_{\mathcal{D}_{T}^{+}}}=\varphi_{\operatorname{KM}}^{E}$, see [5, Equation. (5.10)]. Hence, the two kernels
$\Theta^{V}_{\varphi}\in C^{\infty}\left(\mathscr{H}\right)\otimes\Omega^{q}(Y)$ (3.56)
@@
$\Theta^{T}_{\varphi}\in C^{\infty}\left(\mathscr{H}^{d}\right)\otimes\Omega^{q}(Y_{T})$ (3.57)
@@
$C^{\infty}\left(\mathscr{H}\right)\otimes\Omega^{q}(Y_{T}).$ (3.58)
@@
$\iota^{\ast}\Theta^{T}_{\varphi}=\phi^{\ast}\Theta^{V}_{\varphi}.$ (3.59)
@@
$\Theta^{T}_{\varphi}(\tau,\cdots,\tau,\chi)=\int_{Y_{\chi}}\Theta^{T}_{\varphi}(\tau,\cdots,\tau)=\int_{Y_{\chi}}\phi^{\ast}\Theta^{V}_{\varphi}(\tau)=\int_{\phi_{\ast}Y_{\chi}}\Theta^{V}_{\varphi}(\tau),$ (3.62)
@@
$\Theta^{T}_{\varphi}(\tau,\cdots,\tau,\chi)=\int_{C_{\chi}}\Theta^{V}_{\varphi}(\tau).$ (3.63)
@@
For now we have only considered the case where $E$ is an extension of a field $F$, but we want to consider more general étale algebras of the form $E=E_{1}\times\cdots\times E_{r}$, together with an involution fixing $F=F_{1}\times\cdots\times F_{r}$. Here each of the $F_{i}$ is a totally real field of degree $d_{i}$, that is the fixed field of $E_{i}$. Suppose that $(V,Q)$ is the restriction of scalars of $(E,Q_{\alpha})$ from $F$ to $\mathbb{Q}$, as in 2.1. Then $V$ is a vector space of dimension $2d$ where $d=d_{1}+\cdots+d_{r}$. Furthermore, we can split the quadratic space as $V=V_{1}\oplus\cdots\oplus V_{r}$, where $V_{i}$ is the restriction of scalars of $E_{i}$ from the field $F_{i}$ to $\mathbb{Q}$. We can restrict the embedding $\operatorname{SO}_{V}(\mathbb{Q})\hookrightarrow\operatorname{Sp}_{4d}(\mathbb{Q})$ given in 3.4 to an embedding $\operatorname{SO}_{V_{1}}(\mathbb{Q})\times\cdots\times\operatorname{SO}_{V_{r}}(\mathbb{Q})\hookrightarrow\operatorname{Sp}_{4d}(\mathbb{Q})$. Its commutator is $\operatorname{SL}_{2}(\mathbb{Q})^{r}$ and we get a seesaw
$\operatorname{SO}_{V}(\mathbb{Q})$$\operatorname{SL}_{2}(\mathbb{Q})^{r}$$\operatorname{SO}_{V_{1}}(\mathbb{Q})\times\cdots\times\operatorname{SO}_{V_{r}}(\mathbb{Q})$$\operatorname{SL}_{2}(\mathbb{Q}).$}
Combining with previous seesaw, we then get
$\operatorname{SO}_{V}(\mathbb{Q})$$\operatorname{SL}_{2}(F_{1})\times\cdots\times\operatorname{SL}_{2}(F_{r})$$\operatorname{SO}_{E_{1}}(F_{1})\times\cdots\times\operatorname{SO}_{E_{r}}(F_{1})$$\operatorname{SL}_{2}(\mathbb{Q}).$}
Let $\varphi_{f}\in\mathcal{S}(V_{\mathbb{A}_{f}})$ be a finite Schwartz function. Let $\varphi=\varphi^{V}_{\operatorname{KM}}\otimes\varphi_{f}$ and the associated theta lift
$\Theta^{V}_{\varphi}\colon H^{q}(Y,\mathbb{C})\longrightarrow M_{d}(\Gamma^{\prime}_{\mathbb{Q}}).$ (3.64)
We have $\mathcal{S}(V_{\mathbb{A}_{f}})\simeq\mathcal{S}(\mathbb{A}_{E,f})\simeq\mathcal{S}(\mathbb{A}_{E_{1},f})\otimes\cdots\otimes\mathcal{S}(\mathbb{A}_{E_{r},f})$ and so we can write $\varphi_{f}$ as a finite sum
@@
of parallel weight one for a congruence subgroup $\Gamma_{i}^{\prime}$ of $\operatorname{SL}_{2}(F_{i})$. We denote by
@@
$\int_{C_{\chi}}\Theta^{V}_{\varphi}(\tau)=\sum_{\beta}\Theta_{\varphi^{\beta,1}}(\tau^{\Delta},\chi_{1})\cdots\Theta_{\varphi^{\beta,r}}(\tau^{\Delta},\chi_{r}).$ (3.70)